Find the sum of all 3 digit natural numbers which contain at least one  odd digit and at least one even digit


 

Grade:10

1 Answers

SAGAR SINGH - IIT DELHI
14 years ago

Dear aritra,

If 1st digit is odd 50% of no.s can be there in last two places.
if it is odd rest 50% can be there.
till 1 to 8 in first place sum of last two=4*((101*100)/2)=4*5050=20200
if 9 rests at 1st place sum=5050/2=2525
so total=20200+2525=22725
but we have to add hundred's place. so sum of hundred's place is 50*(100+200+300+400+500+600+700+800+900)…
so sum=225000+22725=247725

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi